OpenStudy (anonymous):

Find three consecutive odd intergers such that the four times the largest is one more than nine times the smallest interger,

OpenStudy (anonymous):

consecutive YAY! consecutive means, 3 integers. so x,y,z. lets say x<y<z. 4z = 9x+1 and x+2 = y. y+2 = z. substitute :)

OpenStudy (anonymous):

Let's set up a system of equations. Three consecutive odd integers: \[x ~ and ~y = x + 2 ~ and ~z = y + 2\] Four times the largest is one more than nine times the smallest. \[4*z = 9*x+1\]

OpenStudy (anonymous):

^ i win eastmore :P

OpenStudy (anonymous):

sorry. eashmore i meant.